Can you please help me to solve the following:
sue the porperites of logarithms to evaluate this expression without using a calculator.
2log₈4 – 1/3log₈8

Thank you!

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

= log(16) - log(2)
= log(8), whatever the base is.

Answer by stanbon(75887)

2log8(4) – (1/3)log8(8)
-----
2[log(4)/log(8)] - (1/3)*1
---
Let the base on the logs be "2" to get:
---
= 2[2/3] - (1/3)
---
= 3/3
= 1
========================
Cheers,
Stan H.
============